OpenStudy (anonymous):

Holly has a rectangular garden that measures 12 m wide by 14 m long. She wants to increase the area to 255 m^2 by increasing the width and length by the same amount. What will be the length (longer dimension) of the new garden?

OpenStudy (anonymous):

(12 + x)(14 + x) = 255 x^2 + 26x + 168 = 255 x^2 + 26x - 87 = 0 (x + 29)(x - 3) x = -29, 3 x = 3 (increased by) so new dimensions is 15, 17 length is 17 m